DIRC Operations Biweekly Report
Period: Wednesday 2007/03/13 - Wednesday 2007/03/27
Oncall: Alejandro Perez, Bradley Wogsland
Fast Monitoring: Alejandro Perez, Bradley Wogsland
PEP peak lumi: ~9 x 1033
(previous week: ~9 x 1033)
BaBar integrated lumi: 3.41 fb-1
(previous week: 2.84 fb-1)
Run 6 PEP delivery: 3.43 fb-1
(previous week: ~2.9 fb-1)
Run 6 total delivery: 19.2 fb-1
Run 6 total BaBar recorded: ~19.2 fb-1
including 0 fb-1 off-peak data
PEP activities this week:
3 Day ROD 3/26-3/28.
Foreseen work opportunities (PEP down periods):
Future RODs not yet scheduled.
BaBar activities/issues this week:
3 Day ROD 3/26-3/28 - Backward doors were open Monday.
- Paged 3/17 due to bin 11 in the FX status exceding its
limit. There was under very bad trickle injection. DQM told to call
MCC to tell them that the Background was high and they fixed it.
- Tested repaired DFBs from France. Current spare status:
- DFBs - 9 good, 3 require reprogramming, 6 broken
- DCCs - 4 good
- FEE Crates - 2 good
- HV Crates - 1 good shared with DCH
- HV boards - 11 good
- Paged 3/21 about FX status plots due to 3 noisy
channels: Sector 0, DFB 5, channel 12, Sector 0, DFB 11, channel 0, Sector 6,
DFB 9, channel 0. Short term DQM instructions updated until we masked
them during the ROD.
- On 3/22 Nicolas created an expert page in JAS to see the FX status
plots per ROM and added the corresponding plots to our FM plots script.
- Paged 3/23 for slightly low HV bin in sector 8.
- Paged 3/24 for noisy channels in sector 0.
- On 3/26 (during ROD) a trip in EH shut off our HV and Wiener crates.
Wiener crates came back fine but IOC had to be rebooted to
get HV crate 1 to register. (The other 5 crates came back fine.)
- Backward doors opened for the ROD on 3/26, so we
replaced the faulty sector 8 DFB (# 87) and the sector 5 DFB with the
overlarge timing window (# 5A). New DFBs looked good in calibration.
- Jochen tested all the masked channels on 3/27 to verify they still
needed to be masked and added the 3 new channels: Sector 0, DFB 5, channel 12,
Sector 0, DFB 11, channel 0, Sector 6, DFB 9, channel 0.
- On 3/27 the new sector 8 DFB showed the same 8 channels missing in
logbook for details
- Replace two DFBs in sector 1 with a group of 8 innefficient or
flaky channels (require bridge).
- Replace new DFB in sector 8 with 8 dead channels and figure out
what is killing DFBs in that slot.
Bradley, Jose and Alejandro rotating backup & on-call commissioning duties